6x-5y=-3
-4x+y=-5

Found 2 solutions by TimothyLamb, Alan3354:
Answer by TimothyLamb(4379)

---
6x-5y=-3
-4x+y=-5
---
copy and paste the above linear system in standard form into this matrix-method solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
solution:
x = 2
y = 3
---
system matrix
6 -5
-4 1
---
inverse of system matrix
-0.071428571 -0.35714286
-0.28571429 -0.42857143
---
determinant of system matrix = -14
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php

Answer by Alan3354(69443)

ok.
Now what?